# Multispecies inhomogeneous $t$-PushTASEP from antisymmetric fusion

### v2

@@ 5. ${H_{\text{PushTASEP}}}$ from transfer matrices
$\mathcal{H}=D^{-1}_{\bf m}\sum_{k=0}^{n+1}(-1)^{k-1}\dot{T}^{k}(0)-\left(\sum_{j=1}^{L}\frac{1}{x_{j}}\right)\mathrm{Id},$ (51)
where $D_{\bf m}$ is given in (8). It defines a linear operator on each sector $\mathbb{V}({\bf m})$.
@@
Theorem 3.
The Markov matrix ${H_{\text{PushTASEP}}}$ of the $t$-PushTASEP in (10)–(11) is identified with $\mathcal{H}$ (51) based on the transfer matrices in Section 4. Namely the following equality holds in each sector $\mathbb{V}({\bf m})$:
${H_{\text{PushTASEP}}}=\mathcal{H}.$ (52)
Example 4.
Set $n=2,L=4$ following Example 1 and Example 2. From (50) we have
$\begin{split}\sum_{k=0}^{3}(-1)^{k-1}\dot{T}^{k}(0)|0121\rangle&=\frac{(1-t)^{3}(1+t+t^{2})}{x_{2}}|1021\rangle+\frac{(1-t)^{3}}{x_{3}}|1102\rangle+\frac{(1-t)^{3}t^{2}}{x_{3}}|1201\rangle\\ &+\frac{(1-t)^{3}t}{x_{3}}|2101\rangle+\frac{(1-t)^{3}(1+t+t^{2})}{x_{4}}|1120\rangle+\dot{\mathcal{D}}(0)|0121\rangle,\end{split}$ (53)
where $\dot{\mathcal{D}}(0)=\sum_{k=0}^{3}(-1)^{k-1}\dot{\mathcal{D}}_{k}(0)$. (See Example 2 for the definition of $\mathcal{D}_{k}(z)$.) These vectors belong to the sector $\mathbb{V}({\bf m})$ with multiplicity ${\bf m}=(1,2,1)$. Thus we have $D_{\bf m}=(1-t)^{2}(1-t^{3})$ according to (8). The vector (53) divided by $D_{\bf m}$ reproduces Example 1, where the coincidence of the diagonal terms will be shown in (55).
The rest of this section is devoted to the proof of Theorem 3.
@@ 5.1. Diagonal elements
As a warm-up, we first prove (52) for the diagonal matrix elements, i.e.,
$\langle\boldsymbol{\sigma}|{H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle=\langle\boldsymbol{\sigma}|\mathcal{H}|\boldsymbol{\sigma}\rangle.$ (54)
From (10) we know $\langle\boldsymbol{\sigma}|{H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle=-\sum_{j=1}^{L}\frac{[\sigma_{j}\geq 1]}{x_{j}}$. The RHS is calculated as
@@
=(44)${=}D_{\bf m}^{-1}\sum_{j=1}^{L}\frac{1}{x_{j}}\sum_{k=0}^{n+1}(-1)^{k}e_{k}(u^{(\sigma_{j})}_{0},\ldots,u^{(\sigma_{j})}_{n})-\sum_{j=1}^{L}\frac{1}{x_{j}}$
=(47)${=}\sum_{j=1}^{L}\frac{\delta_{\sigma_{j},0}}{x_{j}}-\sum_{j=1}^{L}\frac{1}{x_{j}}=-\sum_{j=1}^{L}\frac{[\sigma_{j}\geq 1]}{x_{j}},$ (55)
@@ 5.2. Reduced diagram and its depth
From now on, we assume $\boldsymbol{\sigma}^{\prime}\neq\boldsymbol{\sigma}$ and concentrate on the off-diagonal elements $\langle\boldsymbol{\sigma}^{\prime}|{H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle$ and $\langle\boldsymbol{\sigma}^{\prime}|\mathcal{H}|\boldsymbol{\sigma}\rangle$. The former is given, from (10), as
$\langle\boldsymbol{\sigma}^{\prime}|{H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle$ $=\sum_{j=1}^{L}\langle\boldsymbol{\sigma}^{\prime}|{H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle_{j},\quad with$ (56a)
$\langle\boldsymbol{\sigma}^{\prime}|{H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle_{j}$ $=\frac{1}{x_{j}}\prod_{1\leq h\leq n}w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h),$ (56b)
where the factor $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)$ has been defined in (11). On the other hand $\langle\boldsymbol{\sigma}^{\prime}|\mathcal{H}|\boldsymbol{\sigma}\rangle$ is given, from (36b) and (51), as
$\langle\boldsymbol{\sigma}^{\prime}|\mathcal{H}|\boldsymbol{\sigma}\rangle$ $=D_{\bf m}^{-1}\sum_{k=0}^{n+1}(-1)^{k-1}\sum_{j=1}^{L}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}\quad with$ (57a)
$\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}$ $=\frac{1}{x_{j}}\sum_{{\bf a}_{1},\ldots,{\bf a}_{L}\in\mathscr{B}^{k}}S(0)^{{\bf a}_{2},{\bf e}_{\sigma^{\prime}_{1}}}_{{\bf a}_{1},{\bf e}_{\sigma_{1}}}\cdots\dot{S}(0)^{{\bf a}_{j+1},{\bf e}_{\sigma^{\prime}_{j}}}_{{\bf a}_{j},{\bf e}_{\sigma_{j}}}\cdots S(0)^{{\bf a}_{1},{\bf e}_{\sigma^{\prime}_{L}}}_{{\bf a}_{L},{\bf e}_{\sigma_{L}}}.$ (57b)
@@
$\prod_{1\leq h\leq n}w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)=D_{\bf m}^{-1}\sum_{k=0}^{n+1}(-1)^{k-1}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}.$ (58)
@@
Table 1. Special values $S(0)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{c}}$ and $\dot{S}(0)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{c}}$ obtained from (33) relevant to $\dot{T}^{k}(0)$. The symbols $\delta$ and $\varepsilon$ are shorthand for $\delta=\delta^{{\bf a}+{\bf e}_{b}}_{{\bf i}+{\bf e}_{c}}$ and $\varepsilon=(-1)^{a_{0}+\cdots+a_{c-1}+i_{0}+\cdots+i_{b-1}}$, respectively. For the nonzero cases with $c\neq b$, we use the fact $a_{c}=1$ which follows from the constraint ${\bf a}+{\bf e}_{b}={\bf i}+{\bf e}_{c}$. Similarly, the sign factor for the $c=b$ case has been set to $\varepsilon=1$. The second line with $c=b$ case is found to be irrelevant and is therefore omitted.
@@
$\dot{S}(0)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{c}}$ 0 ---- - - $\delta\varepsilon t^{a_{c+1}+\cdots+a_{n}}(1-t)$
We depict $x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}$ as in Figure 2, suppressing all the spectral parameters $z/x_{i}$ as they are set to zero. All the vertical arrows from $\sigma_{i}$ to $\sigma^{\prime}_{i}$ with $\sigma_{i}=\sigma^{\prime}_{i}$, corresponding to “diagonal transitions”, are omitted. Moreover, we perform a cyclic shift such that the site $j$ appears in the leftmost position (this is merely for ease of visualization and not essential), attaching it with $\circ$ to indicate that $\dot{S}(0)$ should be applied there, in contrast to $S(0)$ for other sites. Such a diagram will be referred to as reduced diagram. See (59), where ${\bf a}_{i}\in\mathscr{B}_{k}$, $s_{i}\neq r_{i}\in\{0,\ldots,n\}$ for $0\leq i\leq g$ with some $1\leq g<L$.
@@
The diagram should be understood as representing the sum in (57b), where the $L-g-1$ vertical arrows corresponding to the diagonal transitions are suppressed, but their associated vertex weights should still be accounted for. Since the carriers ${\bf a}_{i}$’s remain unchanged when crossing the omitted vertical arrows, the summation reduces to those over ${\bf a}_{0},\ldots,{\bf a}_{g}$, where ${\bf a}_{i+1}={\bf a}_{i}+{\bf e}_{r_{i}}-{\bf e}_{s_{i}}$ $(i\mod g+1)$.
Lemma 5.
$\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}=0$, unless the reduced diagram (59) for it satisfies the conditions
$\{r_{0},\ldots,r_{g}\}=\{s_{0},\ldots,s_{g}\}=\{h_{0},\ldots,h_{g}\},$ (60a)
$(r_{0},s_{0})=(h_{g},h_{0}),\quad(r_{i},s_{i})=(h_{q_{i}},h_{q_{i}+1})$ (60b)
for some sequence $0\leq h_{0}<\cdots<h_{g}\leq n$ and $10\leq q_{i}\leq g-$ 1$(i=1,\ldots,g)$,
@@
From weight conservation, (59) vanishes unless the condition (i) $\{r_{0},\ldots,r_{g}\}=\{s_{0},\ldots,s_{g}\}$ holds as multisets. From Table 1, it also vanishes unless the additional conditions (ii) $r_{0}>s_{0}$, $r_{1}<s_{1},\ldots,r_{g}<s_{g}$ are satisfied. Conditions (i) and (ii) together are equivalent to (60a) and (60b). ∎
The increasing sequence $(h_{0},\ldots,h_{g})$ appearing in Lemma 5 represents the list of particle types moved during the transition $\boldsymbol{\sigma}\rightarrow\boldsymbol{\sigma}^{\prime}$ induced by $\dot{T}^{k}(0)$. We refer to this sequence as the moved particle types. By the definition, $g=(\text{number of moved particle types})-1$.
Suppose the diagram (59) satisfies (60a) and (60b) for some moved particle types. To ensure weight conservation at every vertex, the capacity $k$ of the carriers must be at least a certain value. We define the minimum possible capacity as the depth $d$ of the reduced diagram or the transition $\boldsymbol{\sigma}\rightarrow\boldsymbol{\sigma}^{\prime}$. Clearly, the depth is unaffected by the diagonal part of the transition which is suppressed in the reduced diagram. We refer to the carries whose capacity equals the depth as minimal carries.
Example 6.
@@
Here we have employed the tableau representation (16) for the carriers. The comparison between (b) and (c) demonstrates that the depth depends on the ordering of the vertical arrows $s_{i}\rightarrow r_{i}$, even when they correspond to the same moved particle types.
Example 6 also demonstrates that $d\leq g$ in general, and the union of tableau letters contained in the minimal carriers ${\bf a}_{0},\ldots,{\bf a}_{g}$ coincide with the moved particle types $\{h_{0},\ldots,h_{g}\}$ as sets. Moreover, they are uniquely determined from $\boldsymbol{\sigma}$ and $\boldsymbol{\sigma}^{\prime}$, reducing the sum (59) into a single term. In fact, in the reduced diagram (59), ${\bf a}_{0},\ldots,{\bf a}_{g}\in\mathscr{B}^{k=d}$ are determined by the recursion relation ${\bf a}_{i+1}={\bf a}_{i}+{\bf e}_{r_{i}}-{\bf e}_{s_{i}}$ $(i\mod g+1)$ and the “initial condition”:
${\bf a}_{0}$ =$=\{s_{0}\}\cup\mathscr{S}_{1}\cup\cdots\cup\mathscr{S}_{g},\qquad\mathscr{S}_{i}=\begin{cases}\varnothing&\text{if }\;s_{i}\in\{r_{0},\dots,r_{i-1}\},\\ \{s_{i}\}&\text{otherwise}.\end{cases}$ (62)
To summarize the argument thus far, we have reduced the equality (58) slightly to
$\prod_{1\leq h\leq n}w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)=D_{\bf m}^{-1}\sum_{k=d}^{n+1}(-1)^{k-1}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j},$ (63)
where the lower bound of the sum over $k$ has been increased to the depth $d$ of the transition $\boldsymbol{\sigma}\rightarrow\boldsymbol{\sigma}^{\prime}$. The LHS is either zero or a nonzero rational function of $t$, whereas the RHS involves summations over $k$ as well as over carriers from $\mathscr{B}^{k}$ entering the definition of $x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}$ in (59).
In the following, we divide the proof of (63) into two cases, depending on whether its LHS is nonzero or zero. The RHS in these corresponding situations will be referred to as wanted terms and unwanted terms, respectively. From the definition of $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)$ in Section 2, unwanted terms correspond to the situation $s_{0}\neq 0$. In Example 6, (a) is unwanted while (b) and (c) are wanted.

### v3

@@ 5. $H_{\text{PushTASEP}}$} from transfer matrices
$\mathcal{H}=D^{-1}_{\bf m}\sum_{k=0}^{n+1}(-1)^{k-1}\dot{T}^{k}(0)-\left(\sum_{j=1}^{L}\frac{1}{x_{j}}\right)\mathrm{Id},$ (5.1)
where $D_{\bf m}$ is given in (2.4). It defines a linear operator on each sector $\mathbb{V}({\bf m})$.
@@
Theorem 5.1.
The Markov matrix $H_{\text{PushTASEP}}$} of the $t$-PushTASEP in (2.6)–(2.7) is identified with $\mathcal{H}$ (5.1) based on the transfer matrices in Section 4. Namely the following equality holds in each sector $\mathbb{V}({\bf m})$:
${H_{\text{PushTASEP}}}=\mathcal{H}.$ (5.2)
Example 5.2.
Set $n=2,L=4$ following Example 2.1 and Example 4.1. From (4.16) we have
$\begin{split}\sum_{k=0}^{3}(-1)^{k-1}\dot{T}^{k}(0)|0121\rangle&=\frac{(1-t)^{3}(1+t+t^{2})}{x_{2}}|1021\rangle+\frac{(1-t)^{3}}{x_{3}}|1102\rangle+\frac{(1-t)^{3}t^{2}}{x_{3}}|1201\rangle\\ &+\frac{(1-t)^{3}t}{x_{3}}|2101\rangle+\frac{(1-t)^{3}(1+t+t^{2})}{x_{4}}|1120\rangle+\dot{\mathcal{D}}(0)|0121\rangle,\end{split}$ (5.3)
where $\dot{\mathcal{D}}(0)=\sum_{k=0}^{3}(-1)^{k-1}\dot{\mathcal{D}}_{k}(0)$. (See Example 4.1 for the definition of $\mathcal{D}_{k}(z)$.) These vectors belong to the sector $\mathbb{V}({\bf m})$ with multiplicity ${\bf m}=(1,2,1)$. Thus we have $D_{\bf m}=(1-t)^{2}(1-t^{3})$ according to (2.4). The vector (5.3) divided by $D_{\bf m}$ reproduces Example 2.1, where the coincidence of the diagonal terms will be shown in (5.5).
The rest of this section is devoted to the proof of Theorem 5.1.
@@ 5.1. Diagonal elements
As a warm-up, we first prove (5.2) for the diagonal matrix elements, i.e.,
$\langle\boldsymbol{\sigma}|{H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle=\langle\boldsymbol{\sigma}|\mathcal{H}|\boldsymbol{\sigma}\rangle.$ (5.4)
From (2.6) we know $\langle\boldsymbol{\sigma}|{H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle=-\sum_{j=1}^{L}\frac{[\sigma_{j}\geq 1]}{x_{j}}$. The RHS is calculated as
@@
[…]
@@ 5.2. Reduced diagram and its depth
From now on, we assume $\boldsymbol{\sigma}^{\prime}\neq\boldsymbol{\sigma}$ and concentrate on the off-diagonal elements $\langle\boldsymbol{\sigma}^{\prime}|{H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle$ and $\langle\boldsymbol{\sigma}^{\prime}|\mathcal{H}|\boldsymbol{\sigma}\rangle$. The former is given, from (2.6), as
$\langle\boldsymbol{\sigma}^{\prime}|{H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle$ $=\sum_{j=1}^{L}\langle\boldsymbol{\sigma}^{\prime}|{H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle_{j},\quad{\text{with}}$ (5.6a)
$\langle\boldsymbol{\sigma}^{\prime}|{H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle_{j}$ $=\frac{1}{x_{j}}\prod_{1\leq h\leq n}w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h),$ (5.6b)
where the factor $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)$ has been defined in (2.7). On the other hand $\langle\boldsymbol{\sigma}^{\prime}|\mathcal{H}|\boldsymbol{\sigma}\rangle$ is given, from (4.2b) and (5.1), as
$\langle\boldsymbol{\sigma}^{\prime}|\mathcal{H}|\boldsymbol{\sigma}\rangle$ $=D_{\bf m}^{-1}\sum_{k=0}^{n+1}(-1)^{k-1}\sum_{j=1}^{L}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}\quad{\text{with}}$ (5.7a)
$\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}$ $=\frac{1}{x_{j}}\sum_{{\bf a}_{1},\ldots,{\bf a}_{L}\in\mathscr{B}^{k}}S(0)^{{\bf a}_{2},{\bf e}_{\sigma^{\prime}_{1}}}_{{\bf a}_{1},{\bf e}_{\sigma_{1}}}\cdots\dot{S}(0)^{{\bf a}_{j+1},{\bf e}_{\sigma^{\prime}_{j}}}_{{\bf a}_{j},{\bf e}_{\sigma_{j}}}\cdots S(0)^{{\bf a}_{1},{\bf e}_{\sigma^{\prime}_{L}}}_{{\bf a}_{L},{\bf e}_{\sigma_{L}}}.$ (5.7b)
@@
$\prod_{1\leq h\leq n}w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)=D_{\bf m}^{-1}\sum_{k=0}^{n+1}(-1)^{k-1}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}.$ (5.8)
@@
Table 1. Special values $S(0)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{c}}$ and $\dot{S}(0)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{c}}$ obtained from (3.12) relevant to $\dot{T}^{k}(0)$. The symbols $\delta$ and $\varepsilon$ are shorthand for $\delta=\delta^{{\bf a}+{\bf e}_{b}}_{{\bf i}+{\bf e}_{c}}$ and $\varepsilon=(-1)^{a_{0}+\cdots+a_{c-1}+i_{0}+\cdots+i_{b-1}}$, respectively. For the nonzero cases with $c\neq b$, we use the fact $a_{c}=1$ which follows from the constraint ${\bf a}+{\bf e}_{b}={\bf i}+{\bf e}_{c}$. Similarly, the sign factor for the $c=b$ case has been set to $\varepsilon=1$. The second line with $c=b$ case is found to be irrelevant and is therefore omitted.
@@
$\dot{S}(0)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{c}}$ 0 --- $\delta\varepsilon t^{a_{c+1}+\cdots+a_{n}}(1-t)$
We depict $x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}$ as in Figure 5, suppressing all the spectral parameters $z/x_{i}$ as they are set to zero. All the vertical arrows from $\sigma_{i}$ to $\sigma^{\prime}_{i}$ with $\sigma_{i}=\sigma^{\prime}_{i}$, corresponding to “diagonal transitions”, are omitted. Moreover, we perform a cyclic shift such that the site $j$ appears in the leftmost position (this is merely for ease of visualization and not essential), attaching it with $\circ$ to indicate that $\dot{S}(0)$ should be applied there, in contrast to $S(0)$ for other sites. Such a diagram will be referred to as reduced diagram. See (5.9), where ${\bf a}_{i}\in\mathscr{B}^{k}$, $s_{i}\neq r_{i}\in\{0,\ldots,n\}$ for $0\leq i\leq g$ with some $1\leq g<L$.
@@
The diagram should be understood as representing the sum in (5.7b), where the Lg1L-g-1 vertical arrows corresponding to the diagonal transitions are suppressed, but their associated vertex weights should still be accounted for. Since the carriers ${\bf a}_{i}$’s remain unchanged when crossing the omitted vertical arrows, the summation reduces to those over ${\bf a}_{0},\ldots,{\bf a}_{g}$, where ${\bf a}_{i+1}={\bf a}_{i}+{\bf e}_{r_{i}}-{\bf e}_{s_{i}}$ $(i\mod g+1)$.
Lemma 5.3.
$\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}=0$, unless the reduced diagram (5.9) for it satisfies the conditions
$\{r_{0},\ldots,r_{g}\}=\{s_{0},\ldots,s_{g}\}=\{h_{0},\ldots,h_{g}\},$ (5.10a)
$(r_{0},s_{0})=(h_{g},h_{0}),\quad(r_{i},s_{i})=(h_{q_{i}},h_{q_{i}+1})$ (5.10b)
for some sequence $0\leq h_{0}<\cdots<h_{g}\leq n$ and $0\leq q_{i}\leq g-1$$(i=1,\ldots,g)$.
@@
From weight conservation, (5.9) vanishes unless the condition (i) $\{r_{0},\ldots,r_{g}\}=\{s_{0},\ldots,s_{g}\}$ holds as multisets. From Table 1, it also vanishes unless the additional conditions (ii) $r_{0}>s_{0}$, $r_{1}<s_{1},\ldots,r_{g}<s_{g}$ are satisfied. Conditions (i) and (ii) together are equivalent to (5.10a) and (5.10b). ∎
Example 5.4.
A reduced diagram (5.9) corresponding to $(h_{0},\ldots,h_{3})=(0,2,3,4)$ with $g=3$ and $(r_{0},s_{0}),\ldots,\allowbreak(r_{3},s_{3})=(4,0),(2,3),(0,2),(3,4)$ is shown in (b) in Example 5.5. It is associated with the following motion of particles.
00332244 (5.11)
The increasing sequence $(h_{0},\ldots,h_{g})$ appearing in Lemma 5.3 represents the list of particle types moved during the transition $\boldsymbol{\sigma}\rightarrow\boldsymbol{\sigma}^{\prime}$ induced by $\dot{T}^{k}(0)$. We refer to this sequence as the moved particle types. By the definition,
g=(number of moved particle types)1.g=($\text${number of moved particle types})-1.
Suppose the diagram (5.9) satisfies (5.10a) and (5.10b) for some moved particle types. To ensure weight conservation at every vertex, the capacity $k$ of the carriers must be at least a certain value. We define the minimum possible capacity as the depth $d$ of the reduced diagram or the transition $\boldsymbol{\sigma}\rightarrow\boldsymbol{\sigma}^{\prime}$. Clearly, the depth is unaffected by the diagonal part of the transition which is suppressed in the reduced diagram. We refer to the carries whose capacity equals the depth as minimal carries.
Example 5.5.
@@
[…]
Here we have employed the tableau representation (3.4) for the carriers. The comparison between (b) and (c) demonstrates that the depth depends on the ordering of the vertical arrows $s_{i}\rightarrow r_{i}$, even when they correspond to the same moved particle types.
Example 5.5 also demonstrates that $d\leq g$ in general, and the union of tableau letters contained in the minimal carriers ${\bf a}_{0},\ldots,{\bf a}_{g}$ coincide with the moved particle types $\{h_{0},\ldots,h_{g}\}$ as sets. Moreover, they are uniquely determined from $\boldsymbol{\sigma}$ and $\boldsymbol{\sigma}^{\prime}$, reducing the sum (5.9) into a single term. In fact, in the reduced diagram (5.9), ${\bf a}_{0},\ldots,{\bf a}_{g}\in\mathscr{B}^{k=d}$ are determined by the recursion relation ${\bf a}_{i+1}={\bf a}_{i}+{\bf e}_{r_{i}}-{\bf e}_{s_{i}}$ $(i\mod g+1)$ and the “initial condition”:
${\bf a}_{0}$ $=\{s_{0}\}\cup\mathscr{S}_{1}\cup\cdots\cup\mathscr{S}_{g},\qquad\mathscr{S}_{i}=\begin{cases}\varnothing&\text{if }\;s_{i}\in\{r_{0},\dots,r_{i-1}\},\\ \{s_{i}\}&\text{otherwise}.\end{cases}$ (5.14)
To summarize the argument thus far, we have reduced the equality (5.8) slightly to
$\prod_{1\leq h\leq n}w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)=D_{\bf m}^{-1}\sum_{k=d}^{n+1}(-1)^{k-1}x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j},$ (5.15)
where the lower bound of the sum over $k$ has been increased to the depth $d$ of the transition $\boldsymbol{\sigma}\rightarrow\boldsymbol{\sigma}^{\prime}$. The LHS is either zero or a nonzero rational function of $t$, whereas the RHS involves summations over $k$ as well as over carriers from $\mathscr{B}^{k}$ entering the definition of $x_{j}\langle\boldsymbol{\sigma}^{\prime}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle_{j}$ in (5.9).
In the following, we divide the proof of (5.15) into two cases, depending on whether its LHS is nonzero or zero. The RHS in these corresponding situations will be referred to as wanted terms and unwanted terms, respectively. From the definition of $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)$ in Section 2, unwanted terms correspond to the situation $s_{0}\neq 0$. In Example 5.5, (a) is unwanted while (b) and (c) are wanted.